# Universidad de Alcalá, Alcalá de Henares 28805, Spain

### v1

@@ 4 Discussions and Conclusion
In this work, we presented detailed analyses of rare bidirectional SEP events with two clear velocity dispersions during the onset phase, observed by Solar Orbiter on 2023 July 13 and 2024 July 23, respectively. Both events exhibit two distinct velocity dispersion signatures with opposite anisotropies during their onsets, enabling unique constraints on the path lengths and release times of sunward and anti-sunward energetic particles. Based on observed differences in release times, path lengths, particle intensities, and spectral indices between the sunward and anti-sunward streaming populations, together with in-situ identification of SFRs, we suggest that the anti-sunward streaming particles may originate primarily from solar flare acceleration, while the sunward streaming population could be associated with shock acceleration. In typical SEP events, flare-accelerated and shock-accelerated particles often propagate along the same interplanetary magnetic field lines to the observer, making it challenging to distinguish their contributions. However, for a spacecraft located within SFRs, the unique magnetic configuration provides an opportunity to separate the two sources because of their different path lengths.
@@
In turn, estimated path lengths for sunward- and anti-sunward-streaming particles provide a valuable means to constrain the topology of magnetic flux ropes. In both events, the presence of two possible distinct particle sources suggests a scenario in which both legs of the SFR remain magnetically connected to the Sun. Consequently, the central axial length of the SFR may be inferred by combining the measured particle path lengths with the estimated magnetic helicity of the flux rope, as discussed in Appendix B. However, in-situ observations from a single spacecraft limit the ability to unambiguously determine the global structure of the SFR. Coordinated multi-spacecraft observations are required to place more stringent constraints on the topology of these structures.
@@
Although the two events share a similar potential explanation, notable differences in local observational features are evident. Event 1 exhibits clear bidirectional particle flows with minimal scattering, leading to strong anisotropies and distinct velocity dispersion signatures for both sunward and anti-sunward particle populations. Moreover, the SFR structure associated with the onset of event 1 is identified by the PyGS algorithm and supported by bidirectional suprathermal electrons. In contrast, event 2 occurs within a more turbulent magnetic environment. Consequently, the SFR in event 2 is not reliably detected by PyGS and can only be inferred from wavelet analysis of magnetic field. Additionally, bidirectional electron signatures in event 2 are less pronounced compared to event 1. Given the possible explanations discussed above, the associated SFRs in both events likely originated from the Sun and are probably linked to small CMEs. Although their spatial scales are smaller compared to large-scale flux ropes observed within MCs, the underlying magnetic topology may be similar. In both events, the particle propagation is governed by the closed nature of the flux rope, enabling sunward-streaming SEPs to be observed at Solar Orbiter. The identification of two velocity dispersions within the SFR intervals further supports a flux rope geometry that guides particles over extended distances along curved magnetic field lines, analogous to those in large MC events (e.g., Leske et al., 2012; Gómez-Herrero et al., 2017). The frequent occurrence of SFRs (Hu et al., 2018; Chen & Hu, 2022), some of which may originate from the Sun, suggests that a significant number of SEP events might be influenced by these elusive structures, which play a role in modulating particle transport.
@@
Bidirectional SEP events with distinct velocity dispersion can also be explained by the mirror effect. For example, Li et al. (2020) reported an SEP event at where two electron populations were observed streaming away from and toward the Sun with similar release times but different path lengths. They suggested that this was due to reflection occurring beyond , rather than transport along a closed magnetic field topology. Under the mirror scenario, the sunward and anti-sunward particle populations are expected to exhibit comparable release times, particle intensities, and spectral indices. In contrast, our study finds large discrepancies in intensity and spectral slope between the two populations, with the sunward source significantly stronger, making the mirroring explanation unlikely in this case. This also indicates that the clear detection of bidirectional SEP events associated with flux ropes requires more specific conditions. Most importantly, in a scattering environment, the first injection must be much weaker than the second. Otherwise, the initial particle population may dominate and mask the velocity dispersion and anisotropy signatures of the second injection. Such intensity differences might be more significant at higher energies due to varying acceleration efficiencies at different sites, which might explain why sunward SEPs are primarily observed in the HET energy channels.
@@
In summary, our observations demonstrate that SFRs might be involved in shaping the transport and distribution of solar energetic particles in the inner heliosphere. The bidirectional SEP events presented here provide evidence for the possible interplay between injections of different source particles and transient magnetic structures. Continued multi-spacecraft and high-resolution observations, combined with improved modelling efforts, will be essential to further unravel the topology of magnetic flux ropes and the particle transport process in bidirectional SEP events.
@@ Acknowledgements.
- Rodríguez-García et al. (2025) Rodríguez-García, L., Gómez-Herrero, R., Dresing, N., et al. 2025, A&A, 694, A64
@@
- Rodríguez-Pacheco et al. (2003) Rodríguez-Pacheco, J., Cid, C., Blanco, J., & Sequeiros, J. 2003, Solar Physics, 213, 121
- Rodríguez-Pacheco et al. (2020) Rodríguez-Pacheco, J., Wimmer-Schweingruber, R. F., Mason, G. M., et al. 2020, A&A, 642, A7
@@ A.1 Directional electron measurements from EPT
Figure 6 presents the color-coded electron intensities measured by the sunward, anti-sunward, north, and south telescopes of SolO/EPT for the 2023 July 13 and 2024 July 23 events. A comparison of the intensities among different telescopes reveals a pronounced anisotropy in the electron distribution. Analysis of the first-order anisotropy indicates the presence of bidirectional electron flows, with a majority of electrons streaming toward the Sun. The intensity of anti-sunward electrons is less pronounced, likely due to high background levels, which limit the ability to perform precise VDA.
@@ A.2 Identification of magnetic flux ropes
where () represents the forward (backward) propagating modes with respect to the mean magnetic field orientation, and and represent the energy density in forward and backward propagating modes, respectively. Alfvénic fluctuations are characterized by a high cross helicity (), indicating dominant energy in either the or modes, and a low residual energy (), which reflects a balance between kinetic and magnetic energies. In contrast, magnetic energy typically dominates over kinetic energy in SFRs, leading to a negative .
Fig. 7 shows a zoom-in view of wavelet spectrograms for the events on July 13, 2023, and July 23, 2024. The period shown corresponds only to the onset phase of the SEP events. The top two panels display the magnetic field and velocity components. The bottom three panels present the wavelet spectrograms of the normalized magnetic helicity , the normalized cross helicity , and the normalized residual energy . We choose selection criteria ( and ) to identify the possible SFRs, indicated by the black contours. These structures are related to high values of , which indicates that the structures are Alfvénic. For the July 13, 2023 event, the largest flux rope displayed in the left panels persists from approximately 05:00 to 08:00 UT. For the July 23, 2024 event, the largest flux rope shown in the right panels lasts from about 00:00 to 05:00 UT. Taken together, the selection criteria and high values suggest that these structures might be Alfvénic SFRs.
Appendix B Estimation of central axis length of magnetic flux ropes
Figure 8 shows the Grad-Shafranov reconstruction of the flux rope observed by Solar Orbiter between 2023 July 13, 03:21:36 and 08:45:44 UT, which is related to the onset of the bidirectional SEPs. The color scale represents the axial magnetic field component (), while the white arrows indicate the direction and relative strength of the transverse field components (). The reconstructed structure displays a well-defined helical magnetic topology, with a central core of enhanced and rotation of about the axis.
To constrain the global geometry of the flux rope, particularly its central axial length, we combine the path length inferred from VDA with the helical configuration revealed by GS reconstruction. Although the degree of helical winding may vary across the flux rope, we approximate the overall twist by using the mean value of the ratio along the spacecraft trajectory. This ratio quantifies the average helical winding of magnetic field lines about the central axis. We further assume that the effect of axis curvature on the total path length can be neglected, and the magnetic field exhibits approximate cylindrical symmetry. The helical path length is then estimated as

### v2

@@ 4 Discussions and conclusion
In this work we have presented detailed analyses of rare bidirectional SEP events with two clear velocity dispersions during the onset phase, observed by Solar Orbiter on July 13, 2023, and July 23, 2024. Both events exhibit two distinct velocity dispersion signatures with opposite anisotropies during their onsets, enabling unique constraints on the path lengths and release times of sunward and anti-sunward energetic particles. Based on observed differences in release times, path lengths, particle intensities, and spectral indices between the sunward and anti-sunward streaming populations, together with in situ identification of SFRs, we suggest that the anti-sunward streaming particles may originate primarily from solar flare acceleration, while the sunward streaming population could be associated with shock acceleration. In typical SEP events, flare-accelerated and shock-accelerated particles often propagate along the same interplanetary magnetic field lines to the observer, making it challenging to distinguish their contributions. However, for a spacecraft located within SFRs, the unique magnetic configuration provides an opportunity to separate the two sources because of their different path lengths.
@@
In turn, estimated path lengths for sunward- and anti-sunward-streaming particles provide a valuable means to constrain the topology of magnetic flux ropes. In both events, the presence of two possible distinct particle sources suggests a scenario in which both legs of the SFR remain magnetically connected to the Sun. Consequently, the central axial length of the SFR could be inferred by combining the measured particle path lengths with the estimated magnetic helicity of the flux rope, as discussed in Appendix B. However, in situ observations from a single spacecraft limit the ability to unambiguously determine the global structure of the SFR. Coordinated multi-spacecraft observations are required to place more stringent constraints on the topology of these structures.
@@
Although the two events share a similar potential explanation, notable differences in local observational features are evident. Event 1 exhibits clear bidirectional particle flows with minimal scattering, leading to strong anisotropies and distinct velocity dispersion signatures for both sunward and anti-sunward particle populations. Moreover, the SFR structure associated with the onset of event 1 is identified by the PyGS algorithm and supported by bidirectional suprathermal electrons. In contrast, event 2 occurs within a more turbulent magnetic environment. Consequently, the SFR in event 2 is not reliably detected by PyGS and can only be inferred from wavelet analysis of the magnetic field. Additionally, bidirectional electron signatures in event 2 are less pronounced compared to event 1. Given the possible explanations discussed above, the associated SFRs in both events likely originated from the Sun and are probably linked to small CMEs. Although their spatial scales are smaller compared to large-scale flux ropes observed within MCs, the underlying magnetic topology may be similar. In both events, the particle propagation is governed by the closed nature of the flux rope, enabling sunward-streaming SEPs to be observed at Solar Orbiter. The identification of two velocity dispersions within the SFR intervals further supports a flux rope geometry that guides particles over extended distances along curved magnetic field lines, analogous to those in large MC events (e.g. Leske et al., 2012; Gómez-Herrero et al., 2017). The frequent occurrence of SFRs (Hu et al., 2018; Chen & Hu, 2022), some of which may originate from the Sun, suggests that a significant number of SEP events could be influenced by these elusive structures, which play a role in modulating particle transport.
@@
Bidirectional SEP events with distinct velocity dispersion can also be explained by the mirror effect. For example, Li et al. (2020) report an SEP event at where two electron populations were observed streaming away from and towards the Sun with similar release times but different path lengths. They suggest that this was due to reflection occurring beyond , rather than transport along a closed magnetic field topology. Under the mirror scenario, the sunward and anti-sunward particle populations are expected to exhibit comparable release times, particle intensities, and spectral indices. In contrast, our study finds large discrepancies in intensity and spectral slope between the two populations, with the sunward source significantly stronger, making the mirroring explanation unlikely in this case. This also indicates that the clear detection of bidirectional SEP events associated with flux ropes requires more specific conditions. Most importantly, in a scattering environment, the first injection must be much weaker than the second. Otherwise, the initial particle population can dominate and mask the velocity dispersion and anisotropy signatures of the second injection. Such intensity differences might be more significant at higher energies due to varying acceleration efficiencies at different sites, which might explain why sunward SEPs are primarily observed in the HET energy channels.
@@
In summary, our observations demonstrate that SFRs might be involved in shaping the transport and distribution of SEPs in the inner heliosphere. The bidirectional SEP events presented here provide potential evidence of an interplay between injections of different source particles and transient magnetic structures. Continued multi-spacecraft and high-resolution observations, combined with improved modelling efforts, will be essential to further unravel the topology of magnetic flux ropes and the particle transport process in bidirectional SEP events.
@@ Acknowledgements.
- Rodr´ıguez-Garc´ıa et al. (2025) Rodríguez-García, L., Gómez-Herrero, R., Dresing, N., et al. 2025, A&A, 694, A64
@@
- Rodr´ıguez-Pacheco et al. (2003) Rodríguez-Pacheco, J., Cid, C., Blanco, J., & Sequeiros, J. 2003, Solar Physics, 213, 121
- Rodr´ıguez-Pacheco et al. (2020) Rodríguez-Pacheco, J., Wimmer-Schweingruber, R. F., Mason, G. M., et al. 2020, A&A, 642, A7
@@ A.1 Directional electron measurements from EPT
Figure 6 presents the colour-coded electron intensities measured by the sunward, anti-sunward, north, and south telescopes of Solar Orbiter/EPT for events 1 and 2. A comparison of the intensities among different telescopes reveals a pronounced anisotropy in the electron distribution. Analysis of the first-order anisotropy indicates the presence of bidirectional electron flows, with a majority of electrons streaming towards the Sun. The intensity of anti-sunward electrons is less pronounced, likely due to high background levels, which limit the ability to perform precise VDA.
@@ A.2 Identification of magnetic flux ropes
where () represents the forward (backward) propagating modes with respect to the mean magnetic field orientation, and and represent the energy density in forward and backward-propagating modes, respectively. Alfvénic fluctuations are characterized by a high cross helicity (), indicating dominant energy in either the or modes, and a low residual energy (), which reflects a balance between kinetic and magnetic energies. In contrast, magnetic energy typically dominates over kinetic energy in SFRs, leading to a negative .
Figure 7 shows a zoom-in view of wavelet spectrograms for events 1 and 2. The period shown corresponds only to the onset phase of the SEP events. The top two panels display the magnetic field and velocity components. The bottom three panels present the wavelet spectrograms of the normalized magnetic helicity , the normalized cross helicity , and the normalized residual energy . We chose selection criteria ( and ) to identify the possible SFRs, indicated by the black contours. These structures are related to high values of , which indicates that the structures are Alfvénic. For event 1, the largest flux rope displayed in the left panels persists from approximately 05:00 to 08:00 UT. For event 2, the largest flux rope shown in the right panels lasts from about 00:00 to 05:00 UT. Taken together, the selection criteria and high values suggest that these structures might be Alfvénic SFRs.
Appendix B Estimation of the central axis length of magnetic flux ropes
Figure 8 shows the GS reconstruction of the flux rope observed by Solar Orbiter between 03:21:36 and 08:45:44 UT on July 13, 2023, which is related to the onset of the bidirectional SEPs. The colour scale represents the axial magnetic field component (), while the white arrows indicate the direction and relative strength of the transverse field components (). The reconstructed structure displays a well-defined helical magnetic topology, with a central core of enhanced and rotation of about the axis.
To constrain the global geometry of the flux rope, particularly its central axial length, we combined the path length inferred from VDA with the helical configuration revealed by GS reconstruction. Although the degree of helical winding may vary across the flux rope, we approximated the overall twist by using the mean value of the ratio along the spacecraft trajectory. This ratio quantifies the average helical winding of magnetic field lines about the central axis. We further assumed that the effect of axis curvature on the total path length can be neglected, and the magnetic field exhibits approximate cylindrical symmetry. The helical path length is then estimated as